# Evaluating Test-Time Scaling LLMs for Legal Reasoning: OpenAI o1, DeepSeek-R1, and Beyond

### v1

@@ 3.1.2 English Legal Tasks
Legal Argument Reasoning(LAR)  Chlapanis et al. (2024) Acc 200
The English legal reasoning tasks are mainly sourced from LegalBench  Guha et al. (2023), a collaboratively constructed legal reasoning benchmark consisting of 162 tasks covering six different types of legal reasoning. Besides, we also add a new legal argument reasoning task proposed by  Chlapanis et al. (2024). The tasks are listed as follows:
@@
3.2 Models
To evaluate the legal reasoning capabilities of LLMs, we test a range of models developed by different companies, including both general-purpose LLMs and those incorporating test-time scaling.
Table 3: LLMs used for legal reasoning evaluation
Category Model Name Source
General LLMs GPT-4o OpenAI
Llama-3.1-405B Meta
Qwen2.5-72B-Instruct Alibaba
Deekseek-V3 DeepSeek
Test Time Scaling LLMs DeepSeek-R1 DeepSeek
OpenAI-o1-preview OpenAI
OpenAI-o1-mini OpenAI
GLM-zero-preview Zhipu
QwQ-32B-Preview Alibaba
Table 3 summarizes the LLMs employed in our study. For general-purpose LLMs, we employ OpenAI’s GPT-4o, DeepSeek-V3, Meta’s Llama-3.1-405B, and Alibaba’s Qwen2.5-72B-Instruct. Meanwhile, for test-time scaling LLMs, we adopt DeepSeek-R1, OpenAI’s o1-preview, and o1-mini, alongside Zhipu’s GLM-zero-preview and Alibaba’s QwQ-32B-Preview.
3.3 Experiment Setting
During evaluation, we retrieve LLMs responses via API requests. Tailored prompts are designed for each task, ensuring a clear structure for the expected outputs. Additionally, API request parameters are adjusted based on the specific LLMs. Detailed task prompts can be found in Appendix A.
4 Experiment Results
@@
In this section, we detail and analyze the performance of LLMs on legal tasks across both Chinese and English contexts.
4.1 Chinese Legal Task Results
Table 4: The result of Chinese legal task. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
Test Time Scaling LLMs
QwQ-32B-Preview 78.97% 56.00% 73.98% 60.70% 67.04% 87.04% 83.82% 27.75% 56.5% 34.93% 62.00%
As shown in table 4, DeepSeek-R1 demonstrates strong performance in Chinese tasks. However, all LLMs perform poorly on the CFE task, indicating its challenging nature. This may be due to the complex reasoning required, as the CFE task demands that models understand legal cases and infer the correct issue from a set of disputes. Moreover, the legal judgment prediction task for cases with multiple defendants proves more difficult for LLMs than for those with a single defendant. In contrast, the multi-defendant charge prediction task is relatively straightforward for LLMs, with DeepSeek-R1 achieving an accuracy of 91.71%—the highest among all tasks. Furthermore, the results indicate that most LLMs underperform on the AR task, suggesting that these models still struggle with recalling fine-grained legal knowledge. In conclusion, LLMs have difficulty with tasks that require complex reasoning, highlighting a key area for future improvement.
4.2 English Legal Task Results
Table 5: The result of English legal task. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 83.64% 82.41% 80.48% 1.207330476 87.87% 84.30% 81.73%
Llama3.1-405B 90.91% 61.11% 70.89% 7.71700315 80.88% 69.59% 83.84%
Qwen2-72B-Instruct 87.27% 82.41% 70.89% 4.806589399 85.29% 77.19% 77.89%
DeepSeek-V3 90.91% 77.78% 75.00% 2.307866126 83.09% 77.19% 85.00%
Test Time Scaling LLMs
DeepSeek-R1 92.73% 78.70% 77.05% 0.246035591 91.79% 85.28% 88.60%
o1-preview 96.36% 86.11% 79.79% 1.092967177 91.18% 86.98% 86.24%
o1-mini 87.27% 61.11% 66.78% 1.376229363 89.34% 73.53% 66.50%
GLM-zero-preview 83.64% 57.41% 65.41% 7.785547624 90.77% 70.76% 78.50%
QwQ-32B-Preview 78.18% 59.26% 64.73% 3.303448944 71.32% 73.41% 81.00%
As shown in table 5, the LLMs generally perform better on English reasoning tasks compared to Chinese ones. Furthermore, the performance of DeepSeek-R1 is comparable to o1-preview, with each achieving the best score three times. Additionally, the LRC and Sara_E tasks appear to be relatively easy for the LLMs, with the o1-preview achieving 96.36% and 91.18%, respectively. However, the NYSJE task seems to pose a challenge for the LLMs, with the highest accuracy being 80.48%. Moreover, most LLMs models struggle with the Sara_N task, except for DeepSeek-R1. These results suggest that most LLMs may currently lack the necessary capability to handle tasks related to judicial ethics and complex tax calculations.
5 Analysis
In this section, we further explore the performance of LLMs in two key legal reasoning tasks under both Chinese and English settings.
5.1 LLMs Performance in LJP Task
In this section, we evaluate the performance of LLMs on single-defendant and multi-defendant LJP tasks across three datasets: CAIL2018, CMDL, and MultiLJP. The primary difference between these tasks is the number of defendants involved. Additionally, we detail the results for each subtask within these datasets.
Table 6: The CAIL2018 subtask result. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 65.37% 90.67% 77.44% 37.33%
Llama3.1-405B 67.23% 86.00% 83.00% 41.33%
Qwen2-72B-Instruct 76.50% 93.44% 94.44% 50.33%
DeepSeek-V3 77.03% 94.44% 94.78% 50.67%
Test Time Scaling LLMs
DeepSeek-R1 78.00% 95.00% 95.67% 52.00%
o1-preview 76.63% 94.33% 96.67% 48.33%
o1-mini 59.63% 90.33% 62.67% 34.33%
GLM-zero-preview 71.98% 87.11% 93.28% 44.67%
QwQ-32B-Preview 73.98% 90.00% 92.17% 48.33%
Table 7: The CMDL subtask results. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 61.08% 85.57% 81.38% 27.50%
Llama3.1-405B 49.26% 77.86% 58.47% 20.91%
Qwen2-72B-Instruct 64.39% 87.57% 87.15% 29.92%
DeepSeek-V3 63.94% 88.05% 83.69% 31.03%
Test Time Scaling LLMs
DeepSeek-R1 68.48% 92.32% 91.19% 33.58%
o1-preview 66.71% 91.00% 91.29% 30.07%
o1-mini 42.59% 82.22% 30.98% 21.56%
GLM-zero-preview 55.27% 78.97% 73.07% 24.14%
QwQ-32B-Preview 60.70% 79.82% 82.66% 29.90%
Table 8: The MultiLJP subtask results. The best score is bold, and the second-best score is underlined.
[…]
General LLMs
GPT-4o 61.79% 84.72% 90.45% 23.1%
Llama3.1-405B 57.02% 74.13% 81.4% 25.9%
Qwen2-72B-Instruct 61.41% 83.82% 87.78% 24.84%
DeepSeek-V3 61.97% 81.95% 86.61% 28.51%
Test Time Scaling LLMs
DeepSeek-R1 67.15% 85.46% 92.15% 34.65%
o1-preview 63.05% 84.11% 89.14% 27.68%
o1-mini 39.47% 70.07% 36.73% 18.56%
GLM-zero-preview 55.82% 80.03% 68.62% 28.07%
QwQ-32B-Preview 67.04% 85.66% 89.6% 36.17%
From table 6,7 and 8 above, We can observe that the performance of LLMs on the CAIL2018 dataset is superior to that on CMDL and MultiLJP across all subtasks. This may be attributed to the relative simplicity of single-defendant cases, which facilitates comprehension and judgment prediction by the models. However, the o1-mini model performs worst in the article prediction task, with F1 scores of 30.98%, 36.73%, and 62.67% respectively. Additionally, the DeepSeek-R1 model demonstrates superior legal reasoning capabilities compared to the other LLMs in the LJP task.
5.2 LLMs Performance in Scalr Task
In this section, we further analyze the performance of LLMs in the scalr task. This task requires LLMs to comprehend legal language, rather than simply memorizing specific legal knowledge. Similarly to the LegalBench setting, we randomly sample 30% of the full dataset (comprising 172 questions) to serve as our test set. Following the results from LegalBench  [^4] https://hazyresearch.stanford.edu/legalbench/tasks/scalr.html, we compare the performance of some typical LLMs on this task.
Table 9: Performance comparison of different models in scalr task. The best score is bold, and the second-best score is underlined.
[…]
Human 86.00%
DeepSeek-R1 85.28%
o1-preview 86.98%
o1-mini 73.53%
GPT-4o 84.30%
gpt-4-0314 81.40%
gpt-3.5-turbo-0301 64.0%
As shown in table 9, the o1-preview model outperforms human performance, with DeepSeek-R1 ranking second. The latest OpenAI models exceed gpt-3.5-turbo-0301, and our evaluation reveals that recent LLMs also surpass gpt-3.5-turbo-0301, underscoring significant advancements in legal reasoning capabilities.
6 Error Analysis
Based on the experimental results in Section 4, DeepSeek-R1 outperforms other LLMs in most Chinese legal tasks and achieves performance comparable to o1-preview in English tasks. Consequently, we perform an error analysis of R1’s performance on several representative tasks. For Chinese legal tasks, we examine the IAPE, CFE, LJP, and AR tasks; for English legal tasks, we analyze the CPC and NYSJE tasks.
Figure 2: R1 error types across typical tasks.
6.1 IAPE Error Analysis
In the IAPE task, we randomly select 30 incorrect cases and identify two main error types: inconsistent objects and incomplete rebuttals. The inconsistent objects occurs when the model’s prediction doesn’t address the same legal subject as the given argument. The incomplete rebuttals happens when the prediction is not the best choice compared to the correct one. As shown in panel (a) of figure 2, 93% involved inconsistencies in the subject of discussion, while 7% had incomplete rebuttals. After analyzing the thinking process of R1, we conclude the reasons as follows:
(1) In the thinking process of R1, it fails to grasp the core subject in the plaintiff’s argument, making it hard to identify the correct iterative pair.
(2) The R1 fails to identify the argument that directly interacts with the plaintiff, instead selecting a response that comes after a lengthy reasoning process. This excessive reasoning leads to an incorrect answer.
We provide two examples of R1’s flawed thinking in the IAPE task, as shown in table 10.
Table 10: Examples of R1’s flawed thinking in the IAPE task.
Type Example
6.2 CFE Error Analysis
In this task, we also randomly select 30 incorrect cases and category the errors into four types based on the number of disputes incorrectly extracted: complete deviation, major deviation, partial deviation, and minor deviation. In the 30 examples we randomly analyze, 67% exhibit complete deviation, while 13% show minor deviations. The remaining two categories account for 10%. We analyze the thinking process of R1 in error cases, concluding the reasons as follows:
(1) R1 lacks sufficient legal knowledge in certain areas, such as the concept of "an equal division of common property", which involves two key aspects: property division and property nature determination. However, R1 fails to identify the determination of property nature. Similarly, the issue of "arbitration requests for double wages exceeding the statute of limitations" involves two components: whether double wages are due and the statute of limitations for arbitration.
(2) R1 sometimes struggles with understanding more complex cases or the meaning of some disputed issues, which hinders accurate identification of the dispute issue.
(3) Some disputes, such as the determination of loss compensation amount and the determination of payment amount, are confusing to R1, making it difficult to identify the correct one.
The table 11 below provides examples illustrating these reasons.
Table 11: Examples of R1’s flawed thinking in the CFE task.
Type Example
6.3 LJP Error Analysis
In the error analysis of the LJP task, we analyze the sentence prediction task of CAIL2018, CMDL, and MultiLJP. We also randomly select 30 wrong cases and divide the error types into 2 types, underestimation of the sentence and overestimation of the sentence.
As shown in figure2 (c), (d), and (e), in the CAIL2018, 63% of the cases involve an overestimation of the punishment, while 37% reflect an underestimation. In the MultiLJP, we observe that 60% of the incorrect cases involve an overestimation of the punishment, and 40% show an underestimation of the sentence. In the CMDL, 53% of the cases involve an underestimation of the sentence, while 47% reflect an overestimation of the punishment.
We conclude the reasons as follows:
(1) There exist some hallucinations in the thinking process, including highlighting or ignoring the sentencing factors and confusing the crimes. When R1 predicts the wrong charges of cases, the basic sentence is away from the ground truth, resulting in the wrong sentence.
(2) There are some outdated legal provisions in the training data, R1 predicts the sentence based on these articles, resulting the wrong sentence of defendants, such as the articles of operating a gambling house.
(3) We find that R1 predicts the sentence of the current case based on judgment of the similar cases in its training data. Therefore, when it selects the incorrectly similar case, the sentence is away from the ground truth.
Table 12 presents examples illustrating the three reasons.
Table 12: Examples of R1’s flawed thinking in the LJP task.
Type Example
6.4 AR Error Analysis
In the AR task, we also randomly select 30 incorrect cases from R1 and analyze the error types. We summarize four error types: confusion between articles, where the model incorrectly substitutes content from one legal article for another; content fabrication, where R1 generates an entirely fictional article that doesn’t exist in the actual legal corpus; incomplete recitation, where the model omits essential portions of the legal article; and outdated references, where R1 recites previous versions of articles that have since been amended or updated. As shown in figure 2 (f), we observe that in 73% of cases, the model creates confusion between articles, and 17% of cases involve fabricated articles. Besides, 7% of cases involve incomplete recitation, and 3% of cases are outdated references. Table 13 illustrates examples of these error types. observe
Table 13: Examples of R1’s flawed thinking in the AR task.
Error Types Example
6.5 CPC Error Analysis
In the CPC task, we collect all the incorrect cases of this task from R1 and analyze the error types. As shown in figure 1 (g), the R1 model is more likely to misclassify the case labeled as "No" to "Yes" (57%) than the reverse (43%). We further analyze the reasons behind this:
(1) We find that there are factual hallucinations about the content of citation during the thinking process of R1. In addition, when the model is unclear about the details of the case, hallucinations will also occur, resulting in incorrect judgments.
(2) In this task, correctly understanding the citation is crucial for providing an accurate answer. Although R1 has access to the full case details, a deviation in understanding the case also leads to the wrong conclusion.
Table 14 presents examples illustrating the two reasons.
Table 14: Examples of R1’s flawed thinking in the CPC task.
Type Example
6.6 NYSJE Error Analysis
We also collect all the incorrect cases of the NYSJE task from R1 and analyze the error types. As shown in figure 1 (h), we observe that False Positives and False Negatives account for nearly half of the incorrect cases. This task requires LLMs to determine whether a question violates judicial ethics in the New York State Unified Court System. According to the results in the table 5, R1 achieves an accuracy of 77.05%, while the best score is 80.48%, indicating that the task is challenging for R1. We further analyze the reasons behind this.
(1) We find that there are factual hallucinations in ethical guidelines that R1 recalls, which mislead R1 do the wrong conclusion.
(2) In the thinking process of R1, there is an overthinking phenomenon. For example, it adds irrelevant details to the context of the question, which results in incorrect output.
Table 15 provides examples illustrating the two reasons.
Table 15: Examples of R1’s flawed thinking in the NYSJE task.
Type Example
7 Conclusion
This preliminary study evaluates the legal reasoning abilities of LLMs across 18 existing Chinese and English legal reasoning datasets using the latest LLMs models, including the TTS LLMs. We present our initial results and analysis, which provide evidence that the legal reasoning capabilities of TTS LLMs are still underdeveloped. The findings highlight that LLMs continue to struggle with both understanding legal terminology and executing complex legal reasoning. Future development of legal LLMs should prioritize improving these areas to enhance their overall legal reasoning performance.
8 Limitations
In this paper, we evaluate the legal reasoning abilities of LLMs using a collection of Chinese and English legal reasoning tasks. However, the tasks we gathered may not cover the full spectrum of legal reasoning scenarios. In the future, we plan to incorporate additional LLMs and expand the tasks to include subjective legal reasoning tasks. These tasks require applying legal provisions, similar cases, and principles through logical analysis of given facts and rules to reach conclusions. Additionally, defining appropriate evaluation metrics for this task is challenging. Therefore, we will explore and develop suitable metrics to better assess the performance of LLMs in these subjective legal reasoning tasks.
9 Reference
Appendix A Appendix
In this section, we present the instructions for LLMs in the evaluation of the Chinese and English legal tasks.
@@ Appendix A Appendix
Figure 13: The prompt for ArgMine dataset.

### v2

@@ A.2 English Legal Tasks
Legal Argument Reasoning(LAR)  chlapanis-etal-2024-lar Acc 200
The English legal reasoning tasks are mainly sourced from LegalBench  LEGALBENCH, a collaboratively constructed legal reasoning benchmark consisting of 162 tasks covering six different types of legal reasoning. Besides, we also add a new legal argument reasoning task proposed by  chlapanis-etal-2024-lar. The tasks are listed as follows:
@@
Appendix B Appendix B
@@
In this section, we present the instructions provided to LLMs for evaluating legal tasks in both Chinese and English. For details, see Figures 423.
@@
Figure 14: The prompt for IAPE dataset.
@@
Appendix C Appendix C
This section presents examples of flawed reasoning processes observed in several representative tasks.
C.1 IAPE task
Table 10 illustrates typical flawed reasoning identified in the IAPE task.
Table 10: Examples of flawed reasoning in the IAPE task.
Type Example
C.2 CFE task
Table 11 illustrates typical flawed reasoning identified in the CFE task.
Table 11: Examples of flawed reasoning in the CFE task.
Type Example
C.3 LJP task
Table 12 illustrates typical flawed reasoning identified in the LJP task.
Table 12: Examples of flawed reasoning in the LJP task.
Type Example
C.4 AR task
Table 13 illustrates typical flawed reasoning identified in the AR task.
Table 13: Examples of flawed reasoning in the AR task.
Error Types Example
C.5 CPC task
Table 14 illustrates typical flawed reasoning identified in the CPC task.
Table 14: Examples of flawed reasoning in the CPC task.
Type Example
C.6 NYSJE task
Table 15 illustrates typical flawed reasoning identified in the NYSJE task.
Table 15: Examples of flawed reasoning in the NYSJE task.
Type Example
Appendix D Appendix D
D.1 Strategies to Minimize Potential Bias and Improve Data Quality
To minimize potential bias and improve data quality, we adopted several strategies during data collection. First, we incorporated a domain-specific legal terminology dictionary to guide DeepSeek-R1 toward generating legally relevant responses, as prior studies have shown that such dictionaries can significantly enhance terminological accuracy in generation tasks zheng2024finetuninglargelanguagemodels. Second, we designed structured prompt templates with explicit formatting and structural constraints, which helped reduce hallucinations and improve both consistency and legal formality, consistent with recent empirical findings de2024towards. Third, we conducted rigorous human review by randomly sampling each batch of generated data and verifying the reasoning process, with particular attention to the correctness of cited legal articles and the professionalism of the language. In summary, while we acknowledge the inherent risk of bias, these measures have effectively mitigated its impact, keeping it within acceptable limits for the purposes of this study.
D.2 Impact of Retrieval Quality
In Table 3, we observe that high-quality supplementary knowledge and contextual information play a crucial role in improving model performance on knowledge-intensive tasks. To further investigate how retrieval quality influences performance, we conducted a controlled experiment where the only variable was the quality of the retrieved documents.
Specifically, we evaluated model performance under three retrieval settings of different quality levels:
High-quality: Gold-reference passages that directly contain the correct answer (ideal retrieval).
Medium-quality: High-quality content mixed with three unrelated legal articles (moderate noise).
Low-quality: Five randomly selected legal articles unrelated to the input query.
Table 16: Impact of retrieval quality on model performance.
Model CAIL2018 CMDL MultiLJP
DeepSeek-R1-Distill-Qwen-14B 72.03% 50.00% 56.50%
with RAG (High-quality) 74.41% 56.94% 66.28%
with RAG (Medium-quality) 70.90% 49.56% 57.23%
with RAG (Low-quality) 67.40% 45.33% 54.19%
DeepSeek-R1 78.00% 68.48% 67.15%
with RAG (High-quality) 80.10% 72.21% 71.49%
with RAG (Medium-quality) 78.37% 68.23% 67.34%
with RAG (Low-quality) 76.17% 65.95% 64.89%
As shown in Table 16, retrieval quality has a substantial impact regardless of whether the base model(DeepSeek-R1-Distill-Qwen-14B) or DeepSeek-R1 is used. When low-quality or noisy content is retrieved, performance drops significantly—even compared with the setting where no external context is provided.